OpenStudy (anonymous):

The formula for the area of a trapezoid with altitude k and bases m and n is . Solve for m

OpenStudy (anonymous):

A = ((m+n)/2)*k

OpenStudy (anonymous):

1/2(k)(m+n)=A 2A/k=m+n m=2A/k -n

OpenStudy (anonymous):

The formula for the area of a trapezoid with altitude k and bases m and n is \[a=\left(\begin{matrix}k \\ 2\end{matrix}\right) (m+n)\] . Solve for m

OpenStudy (anonymous):

have already done that

OpenStudy (anonymous):

is that matrix notacion ?

OpenStudy (anonymous):

rational expressions